99 Marks: 1 What is the mass (in grams) of the solid product when 6.83 g...

99

Marks: 1

What is the mass (in grams) of the solid product when 6.83 g of iron(II) nitrate reacts in solution with excess sodium hydroxide?

Fe(NO3)2(aq) + NaOH(aq) --> Fe(OH)2(s) + NaNO3(aq) [unbalanced]
Answer #1

Fe(NO3)2 + 2 NaOH= Fe(OH)2 + 2 NaNO3(aq)

Moles of Fe(NO3)2 = mass / molar mass

= 6.83 / 180

= 0.038

Moles of Fe(NO3)2 = moles of Fe(OH)2

= 0.038

Mass of Fe(OH)2 = mole * molar mass

= 0.038 * 90

= 3.42 grams

If you have any query please comment

If you satisfied with the solution please like it thankxx